The composite wall of an oven consists of threematerials, two of which are of known thermal conductivity,kA = 20 W/m•K and kC =50 W/(m•K), and known thickness, LA =0.30m and LC = 0.15m. The third material,B, which is sandwiched between materials A andC, is known thickness, LB = 0.15m, butunknown thermal conductivity kB.



Under steady-state operating conditions, measurementsreveal an outer surface temperature of Ts,o =20ºC, an inner surface temperature of Ts,i= 600ºC, and an oven air temperature ofT¥=800ºC. The inside convection coefficienth is known to be 25 W/(m2•K). What isthe value of kB?
Answer given kB = 1.53W/(m×K)
Answer #1
Concepts and Reason

Plane walls sometimes consist of several layers of different materials. The concept of thermal resistance can be used to determine the rate of steady heat transfer through such ‘composite walls’.

The thermal resistance is analogous to electrical resistance and follows the series and parallel law.

The convective resistance is used at the peripheries of the wall and conductive resistance is used between the walls of similar material.

Fundamentals

The composite wall (consisting of two plane walls in series) and its thermal resistance network for heat transfer are as shown:

Wall 2
Wall 1
h,
ool
T2
h2
T3
k2
Т..
о2
L2
conv,2
R
conv,1
Т.2
Т.
2
o02
3
1
oo1

The formulae for thermal resistances are:

a. Convective resistance at wall 1:

Rconv,1=1Ah1{R_{conv,1}} = \frac{1}{{A{h_1}}}

Here, the area of cross section is A and the convective coefficient at wall 1 is h1{h_1} .

b. Conductive resistance in wall 1:

R1=L1Ak1{R_1} = \frac{{{L_1}}}{{A{k_1}}}

Here, the length of wall 1 is L1{L_1} and its thermal conductivity is k1{k_1} .

c. Conductive resistance in wall 2:

R2=L2Ak2{R_2} = \frac{{{L_2}}}{{A{k_2}}}

Here, the length of wall 2 is L2{L_2} and its thermal conductivity is k2{k_2} .

d. Convective resistance at wall 2:

Rconv,2=1Ah2{R_{conv,2}} = \frac{1}{{A{h_2}}}

Here, the convective coefficient at wall 2 is h2{h_2} .

The rate of heat transfer can be calculated as,

Q˙=T1T2Rconv,1+R1+R2+Rconv,2\dot Q = \frac{{{T_{\infty 1}} - {T_{\infty 2}}}}{{{R_{conv,1}} + {R_1} + {R_2} + {R_{conv,2}}}}

Draw the thermal circuit.

RC
RB
соny,
Ts
Т,
S,o
Т.
S,i

Obtain the thermal resistances per unit area.

Rconv,1=1h{R_{conv,1}} = \frac{1}{h}

Here, the convective coefficient is h.

RA=LAkA{R_A} = \frac{{{L_A}}}{{{k_A}}}

Here, the length and thermal conductivity of wall A are LA{L_A} and kA{k_A} respectively.

RB=LBkB{R_B} = \frac{{{L_B}}}{{{k_B}}}

Here, the length and thermal conductivity of wall B are LB{L_B} and kB{k_B} respectively.

RC=LCkC{R_C} = \frac{{{L_C}}}{{{k_C}}}

Here, the length and thermal conductivity of wall C are LC{L_C} and kC{k_C} respectively.

Calculate the heat flux.

q˙=TTS,iRconv,1\dot q = \frac{{{T_\infty } - {T_{{\rm{S,i}}}}}}{{{R_{conv,1}}}}

Here, the temperature of air is T{T_\infty } and the temperature of the inner wall is TS,i{T_{{\rm{S,i}}}} .

Substitute 800C800^\circ {\rm{C}} for T{T_\infty } , 1h\frac{1}{h} for Rconv,1{R_{conv,1}} , 25W/m2K25{\rm{ W/}}{{\rm{m}}^2} \cdot {\rm{K}} for h, and 600C600^\circ {\rm{C}} for TS,i{T_{{\rm{S,i}}}} .

q˙=8006001/h=8006001/25=5000W/m2\begin{array}{c}\\\dot q = \frac{{800 - 600}}{{1/h}}\\\\ = \frac{{800 - 600}}{{1/25}}\\\\ = 5000{\rm{ W/}}{{\rm{m}}^2}\\\end{array}

Calculate the thermal conductivity of material B.

q˙=TS,iTS,oRA+RB+RC\dot q = \frac{{{T_{{\rm{S,i}}}} - {T_{{\rm{S,o}}}}}}{{{R_A} + {R_B} + {R_C}}}

Here, the outer wall temperature is TS,o{T_{{\rm{S,o}}}} .

Substitute 5000W/m25000{\rm{ W/}}{{\rm{m}}^2} for q˙\dot q , LAkA\frac{{{L_A}}}{{{k_A}}} for RA{R_A} , LBkB\frac{{{L_B}}}{{{k_B}}} for RB{R_B} , LCkC\frac{{{L_C}}}{{{k_C}}} for RC{R_C} , 600C600^\circ {\rm{C}} for TS,i{T_{{\rm{S,i}}}} , 20C20^\circ {\rm{C}} for TS,o{T_{{\rm{S,o}}}} , 20W/mK20{\rm{ W/m}} \cdot {\rm{K}} for kA{k_A} , 50W/mK50{\rm{ W/m}} \cdot {\rm{K}} for kC{k_C} , 0.3 m for LA{L_A} , 0.15 m for LB{L_B} , and 0.15 m for LC{L_C} .

5000=60020LAkA+LBkB+LCkC0.320+0.15kB+0.1550=0.1160.15kB=0.098kB=1.53W/mK\begin{array}{l}\\5000 = \frac{{600 - 20}}{{\frac{{{L_A}}}{{{k_A}}} + \frac{{{L_B}}}{{{k_B}}} + \frac{{{L_C}}}{{{k_C}}}}}\\\\\frac{{0.3}}{{20}} + \frac{{0.15}}{{{k_B}}} + \frac{{0.15}}{{50}} = 0.116\\\\\frac{{0.15}}{{{k_B}}} = 0.098\\\\{k_B} = 1.53{\rm{ W/m}} \cdot {\rm{K}}\\\end{array}

Ans:

The thermal conductivity of material B, kB{k_B} is 1.53W/mK1.53{\rm{ W/m}} \cdot {\rm{K}} .